what is the answer to 1/3x+3=3/4y?
Answer by stanbon(75887)

what is the answer to 1/3x+3=3/4y?
----
Multiply thru by 12 to get:
---
4x+36 = 9y
y = (4/9)x + 4
======================
Cheers,
Stan H.
===========